OpenStudy (anonymous):

-x-y=8 x-3y=8 solve by elimination

OpenStudy (lgbasallote):

add the two -4y = 16 y= -4 x = x = -4

OpenStudy (anonymous):

-4y = 16/y = -4/x = -4